OpenStudy (anonymous):

Find the geometric mean of 99 and 11

OpenStudy (anonymous):

geometric mean of 2 numbers a and b is\[\sqrt{ab}\]

OpenStudy (anonymous):

so geometric mean of 99 and 11 becomes\[\sqrt{99\times11}=\sqrt{9\times11\times11}\]

OpenStudy (anonymous):

i think u can do it from here :)

OpenStudy (anonymous):

Mhmm thanks(:

OpenStudy (anonymous):

welcome :)